### v1

# IMPROVE: Impact of Mobile Phones on Remote Online Virtual Education

Abstract
This work presents the IMPROVE dataset, designed to evaluate the effects of mobile phone usage on learners during online education. The dataset not only assesses academic performance and subjective learner feedback but also captures biometric, behavioral, and physiological signals, providing a comprehensive analysis of the impact of mobile phone use on learning. Multimodal data were collected from 120 learners in three groups with different phone interaction levels. A setup involving 16 sensors was implemented to collect data that have proven to be effective indicators for understanding learner behavior and cognition, including electroencephalography waves, videos, eye tracker, etc. The dataset includes metadata from the processed videos like face bounding boxes, facial landmarks, and Euler angles for head pose estimation. In addition, learner performance data and self-reported forms are included. Phone usage events were labeled, covering both supervisor-triggered and uncontrolled events. A semi-manual re-labeling system, using head pose and eye tracker data, is proposed to improve labeling accuracy. Technical validation confirmed signal quality, with statistical analyses revealing biometric changes during phone use.
Background & Summary
Education is a fundamental societal pillar and has continually evolved throughout history. In recent years, it has been deeply influenced by the digital revolution, which has moved our society from the physical to the digital world. Evidence of this change in recent decades is the growing significance of e-learning or online education, expected to grow exponentially over the next 20 years [1, 2]. Relevant institutions are adopting this model, especially those offering MOOCs (Massive Open Online Courses), due to their ability to reach a broader learner base [3].
Learners have also been significantly impacted by this digital revolution, altering their habits to incorporate electronic devices into their daily routines such as mobile phones and tablets. Since 2012, studies [4, 5] indicated that between 97%-99% of learners own a mobile phone. Moreover, 95% of learners bring their phones to class every day, 92% use these devices to send text messages during class, and 10% even admit to texting during exams. On average, learners spend about 5 hours per day on their mobile phones, predominantly using messaging services (sending an average of 109 text messages per day) or social networks like WhatsApp and Instagram [6, 7].
Therefore, the mobile phone raises numerous questions: Can it become a distraction to learners’ attention? Or conversely, could the absence of a mobile phone cause greater distraction or anxiety? Considering that attention is defined as the ability to concentrate, specifically, to exert conscious cognitive effort on a particular task or stimulus at any given moment [8], and given that various studies suggest it is easy to lose focus due to short-term distractions such as messages, sounds, etc. [9, 10]; it is valid to consider how mobile phones could impact learner performance. This is not a new idea, and various studies have shown that mobile phone use negatively affects learners’ attention, memory, and academic performance. This is mainly because it increases cognitive load by inducing learners to multitask [10, 5, 11, 12]. This effect becomes more noticeable as the class progresses or in long-duration activities [11]. However, learners’ performance can be influenced by their emotional state. For this reason, nomophobia, defined as the fear, stress, or anxiety caused by being unable to access or use a mobile phone, is an important factor to consider. Various studies have demonstrated how this condition affects memory and attention, ultimately impacting learners’ academic performance [13, 11].
It is interesting to note that, although nomophobia can lead to lower academic performance, the previously mentioned studies have shown that just the presence of a mobile phone also reduces learners’ attention levels, presenting a complex challenge that requires in-depth investigation.
The IMPROVE database was obtained in a realistic online learning environment. A total of 120 learners were monitored, divided into three different groups to understand the effects of mobile phone usage. To our knowledge, IMPROVE dataset is the first to provide a comprehensive multimodal acquisition setup including behavioral [1, 14], biometric, and learning analytics data from learners during an authentic online course that considers mobile phone usage. The database aims to facilitate understanding of the effects of mobile phone usage in online education. Additionally, it provides a robust dataset that can be utilized to develop new technologies to enhance learners’ learning processes or ensure secure certification [14, 15, 16, 17, 18].
We employed 16 sensors, recognized as effective indicators for understanding learner behavior [19, 8, 20, 21], including an electrodermal activity (EDA) sensor, an electroencephalogram (EEG) band, 2 pulse meters, an accelerometer, a gyroscope, 3 RGB cameras, 2 Near-Infrared (NIR) cameras, a mouse, a keyboard, a microphone, a monitor, and an eye tracker (see Fig. 1). These sensors enabled us to monitor biometric and behavioral signals throughout 120 online learning sessions, each lasting approximately 30 minutes. During these sessions, learners interacted with educational content related to HTML, including video lectures, reading documents, and completing tests to assess their knowledge both before and after the session (with a total of 16 questions per session). This comprehensive setup allowed us to track not only how learners engaged with the material but also how external factors, such as mobile phone interruptions, affected their focus and learning outcomes.
In total, the database comprises 2.83 terabytes of data, making it one of the most extensive multimodal datasets in this domain. This includes 145 labeled mobile phone interruptions, providing crucial insights into how mobile phone usage affects learners.
Figure 1: Acquisition setup during the data capture using the edBB[1, 14] platform, illustrating all the sensors and devices employed, along with examples of the captured information.
Methods
This section outlines the experimental methods and materials utilized for data collection. The database was obtained in accordance with the Declaration of Helsinki and was approved by the Ethics Committee of the Universidad Autónoma de Madrid (No. 11/04/2023). Furthermore, the learners were informed about the experimental procedure and data acquisition setup, and they signed a written consent form. An informed consent form was used to inform the learners that the obtained information will be used solely for research purposes. Learners received an incentive for participating.
Database: IMPROVE
The database consists of 120 learners from the School of Engineering at the Universidad Autónoma de Madrid (UAM), who were monitored during a real online learning session focused on HTML lasting between 20 and 35 minutes. The session was part of the MOOC called “Introduction to Development of Web Applications” (WebApp MOOC for short) available on the edXaaahttps://www.edx.org/ platform.
Category Subcategory Percentage Number of Learners Average Age
Degree Program Computer Engineering 32.5% 39 21.77
Telecommunications Engineering 30.0% 36 22.06
Biomedical Engineering 25.0% 30 19.80
Data Science and Engineering 3.3% 4 20.00
Computer Engineering and Mathematics 9.2% 11 20.55
HTML Proficiency None 42.5% 51 20.02
Beginner 35.8% 43 21.95
Intermediate 16.7% 20 21.80
Advanced 5% 6 23.67
Medical Conditions No issues 55
Glasses 41
Contact lenses 19
Heart murmur 2
Myopia without correction 4
Overall Averages Overall Average Age 21.19
Average Age of Male Learners 51.83% 61 21.64
Average Age of Female Learners 49.16% 59 20.73
Table 1: Distribution of Learners by Degree Program, HTML Proficiency Levels, Medical Conditions (percentage are not shown because learners can have more than one condition), and Gender, with Average Ages.
Recruitment.
Learners from the School of Engineering at the Universidad Autónoma de Madrid (UAM) were recruited. The School of Engineering was selected because the five different degrees offered require basic computer science knowledge, which aligns well with the online course. To encourage participation, an economic incentive was provided. A total of 170 learners signed up for the study, from which 120 were selected. An initial test was conducted with 6 learners to verify the protocol and monitoring system. Adjustments were made after the initial test, such as: requiring Group 1 learners to provide more detailed responses to messages (for more information about the groups formed, see the subsection Protocol), providing a more detailed explanation of the session to learners before starting, eliminating an initial form to avoid potential suggestion, and making adjustments to the EEG band and the calibration of the eye tracker. The 6 learners from the initial test did not participate in the final data collection. 4 other learners were discarded due to issues during the session and replaced by other participants. The final 120 learners were divided into three groups of 40. The selection process considered various criteria to ensure balanced groups, including gender (61 male, 59 female), HTML proficiency (none, beginner, intermediate, advanced), degree program, among other factors. The learners’ age range was 18 to 32 years old, with an average age of 21 years old (see Table 1). The learners gave their consent for the publication of the database. Each learner was anonymized with an identifier that cannot be linked to their identity.
Protocol.
The objective was to understand the effects of mobile phone use or absence on learner behavior and academic performance. To this end, three different groups were formed:
  1. 1.
    Group 1: Mobile phone use and possession allowed. The device was placed on the learner’s desk, visible to the learner, with sound and vibration activated.
  2. 2.
    Group 2: Mobile phone possession was allowed, but their use was prohibited. The device was also placed on the learner’s desk, but with the screen facing down, and with sound and vibration activated.
  3. 3.
    Group 3: The mobile phone was confiscated during the whole learning session.
Figure 2: Protocol during the learning session to evaluate the impact of mobile phone use on learners’ academic performance. The top section outlines the session protocol, dividing learners into three groups based on their allowed level of interaction with mobile phones. The bottom section illustrates the activities performed by learners and shows how events and activities were labeled for later analysis. A graph corresponding to a user’s attention signal is shown, with clear event labeling aligned with specific activities.
Three supervisors were present throughout the session. The first supervisor was responsible for explaining the session in detail to the learner, describing the operation of the sensors, placing and calibrating them, and ensuring that the platforms were working correctly. The second supervisor used the LOGGE tool [22] (for more details on this tool, see the "Data, platforms and sensors" subsection) to label the moments when the learner interacted with the mobile phone or when it rang, as well as to label other relevant information for the session. The third supervisor was responsible for sending two messages to the learners during key moments of the learning session of both Group 1 and Group 2, reviewing the information recorded by the second supervisor, noting any possible incidents during the session, and assisting the learner in case of doubts. We will differentiate between two different types of interruptions: interventions (the two controlled messages sent by the supervisor) and casual interactions (uncontrolled messages/calls received by the learner out of the interventions). Therefore, interventions are considered controlled events, while interactions are classified as uncontrolled events. The first intervention was received during an explanatory video of the lesson, and the second during the reading of HTML code, which was part of the learner evaluation (see Fig. 2). Learners in Group 1 were required to respond with a detailed reply. The messages were simple questions such as: "What do you think of the cafeteria of our School of Engineering?" or an image of a popular movie like Harry Potter accompanied by the question: "What do you think of this film series? Would you recommend it?" On average, learners took about 26 seconds to respond to the first message and about 36 seconds to respond to the second message. Learners in Group 2 were not allowed to use their mobile phones, so their phones simply rang and vibrated during the interruptions.
Additionally, learners in Group 1 were given the freedom to respond/send messages or calls to their contacts. Each of these interactions was recorded. On average, the learners in Group 1 chose to interact with their mobile phones approximately 5 times (including controlled messages and uncontrolled messages or calls), while the mobile phone in Group 2 rang/vibrated approximately 7 times.
Tasks
Each learner completed a single learning session on HTML from the WebApp MOOC. The session lasted an average of 30 minutes, with a maximum limit of 40 minutes. The tasks performed during the session were as follows:
Before starting the HTML unit:
  • Calibration and Learner Information: Learners were briefed again on the session’s procedures. Sensors, such as the eye tracker and EEG band, were fitted to the learners and then calibrated. Learners were asked to provide information that could affect the sensors, such as the use of glasses/contact lenses, cardiac conditions, handedness (right or left-handed), etc. The specific medical conditions reported by the learners, including their prevalence, are summarized in Table 1.
  • EDA Test: Before starting the session, an initial EDA test was conducted to understand the learner’s stress levels. Learners were also asked to self-report their current state as "Very stressed, stressed, normal, relaxed, extremely relaxed" to obtain their feedback.
  • Form and Pretest: Additionally, they fill out a form with the following 2 questions, i) What is your current level of stress/anxiety? and ii) What is your current level of distraction? This aims to gather learner feedback before the session begins and to validate it with sensor data. We designed a pretest consisting of 8 multiple-choice questions of low to intermediate difficulty about HTML, with two objectives: i) to assess the learners’ current level of HTML knowledge and ii) to balance the three groups.
HTML Unit:
The HTML unit consisted of a theoretical part where learners were introduced to HTML concepts in web design through explanatory videos and reading web documents, and a second part where learners’ understanding was evaluated. This evaluation involved 8 multiple-choice questions. After finishing, learners were able to review their incorrect answers. In neither part were learners allowed to return to a previous activity, pause the videos, or skip forward/backward in the video. The tasks were ordered as follows (see Fig. 2):
  1. 1.
    Videos: Two introductory videos on interactive web and basic HTML concepts. The first video lasted 20 seconds, and the second one lasted 1 minute and 48 seconds. During the second video, a message was sent to the learners in Groups 1 and 2 after 1 minute, coinciding with the explanation of HTML concepts.
  2. 2.
    Document Reading: Two documents further explaining the videos, covering concepts such as HTML structure, different tags, etc.
  3. 3.
    Test: This test was designed to be compared with the pretest, in order to evaluate whether the students had learned during the learning session compared to their initial knowledge. The questions were similar to those in the pretest. Six multiple-choice questions followed by a phase of reading HTML code to answer two more multiple-choice questions. During the HTML code reading, the second message was sent to learners at the beginning of the reading when comprehension of the code required greater attention. However, learners could resume reading the code after answering the message, unlike the video where they could not rewind.
  4. 4.
    Review Phase: Finally, learners accessed a review phase where they could check and understand the mistakes they made.
Learner Feedback Group 1 Group 2 Group 3
Low Medium High Low Medium High Low Medium High
Anxiety at Start 52.63 39.47 7.89 51.43 40.0 8.57 43.24 51.35 5.41
Distraction at Start 47.37 52.63 0.0 51.43 34.29 14.29 67.57 21.62 10.81
Anxiety During Session 60.53 28.95 10.53 60.0 34.29 5.71 64.86 35.14 0.0
Distraction During Session 44.74 39.47 15.79 42.86 45.71 11.43 72.97 21.62 5.41
Session Difficulty 37.5 55.00 7.5 40.0 42.5 17.5 45.0 47.5 7.5
Performance Level 7.89 50.0 42.11 8.57 60.0 31.43 10.81 43.24 45.95
Phone Distraction 26.32 50.0 23.68 40.0 50.0 10.0 - - -
Stress Due to Phone Inaccessibility - - - 73.68 26.32 0.0 95.0 5.0 0.0
Table 2: Learner Feedback from Forms, Divided into Three Levels (Low, Medium, and High) and Expressed as Percentages.
End of the Session:
  • A final EDA Test was conducted to detect post-session stress levels after the session, and their feedback was collected again.
  • Learners completed a post-session Form, which asked about their stress, distraction, difficulty, and performance during the session. Additionally, learners in Groups 1 and 2 were asked about phone distraction, while Group 2 and 3 were asked about stress from phone deprivation. Table 2 presents the responses collected from the self-reported forms.
  • Learners were asked to fill out a form with the following 6 questions:
    • "What was your level of stress/anxiety during the learning session?"
    • "What was your level of distraction during the learning session?"
    • "How would you evaluate the difficulty of the learning session you just completed?"
    • "How would you rate your performance during the learning session?"
    • "Observations: Please add any comments you wish to make after the data collection."
    The final question varied depending on the group to which the learner belonged:
    • For Groups 1 and 2: "How distracted were you by your mobile phone?"
    • For Group 2 and 3: "How stressed were you by not being able to use your mobile phone?"
Information Type Sensors or Platforms Sampling Rate Features
Video 3 RGB cameras
2 Infrared cameras
1 Depth camera
20 Hz - 30 Hz MP4 files with codec H264
Desktop Video Screen 1 Hz MP4 file with codec H264
Audio Microphone 8000 Hz Uncompressed WAV files
Keystroke Keyboard 12 Hz Keypress and key release events
Mouse Dynamics Mouse 895 Hz Mouse events: Move, press/release, drag and drop and mouse wheel spin
EEG Band 1 Hz Power Spectrum Density of five frequency bands. Level of attention (from 0 to 100) and eyeblink strength
Pulse, Stress, temperature and Inertial 2 SmartWatch: Huawei Watch 2 and FitBit Sense with PPG, Gyroscope, Accelerometer, Electrodermal Activity 100 Hz Timestamps and data from the pulse, EDA sensor (skin conductance), body temperature and the inertial sensors (accelerometer, gyroscope, magnetometer)
Visual attention, Gaze, and Eyeblink 2 Eye Tracking Cameras 120 Hz Gaze origin and point, saccades, fixations, gaze event duration, pupil diameter, eyeblink, data quality
Face position, Head Pose and Facial Landmarks edBB platform 30Hz Facial bounding box information, 468 facial landmarks, Euler angles
Context Data edBB platform, LOGGE Data-M2LADS, edX, learner, Forms NA Learner Information (gender, HTML proficiency level, field of study, health issues, etc.), mobile phone usage events, computer details including name, private IP, public IP, MAC address, operating system (OS), architecture, keyboard language, screen resolution, available memory, total memory, start and finish times of each task session, and test answers, etc.
Table 3: Summary of Sensors and Data Collection from the IMPROVE Database.
Data, Platforms and Sensors
Two different platforms were used to monitor the learners: the edX and edBB platforms [1, 14]. The edX platform is an online MOOC platform used to conduct the learning sessions. It also allows the capture of session metadata from the learners, such as test answers, start and end times of tasks, and more. The edBB platform was used to monitor behavioral and biometric signals that have been shown to contain relevant information for understanding learner learning [1, 14, 21, 8, 19, 23, 24] (see Daza et al. [1] for a video demonstrationbbbhttps://www.youtube.com/watch?v=JbcL2N4YcDM). A multimodal acquisition framework was designed, incorporating 16 sources of information/sensors (see Fig. 1 and Table 3):
  • Cameras: 2 Logitech C170 cameras (side and overhead) operating at 20 Hz with a resolution of $640\times 480$, and one front-facing RealSense camera were used. The RealSensecccwww.intelrealsense.com/wp-content/uploads/2020/06/Intel-RealSense-D400-Series-Datasheet-June-2020.pdf camera contains one RGB camera and two NIR cameras, with dimensions of 90 mm length $\times$ 25 mm depth $\times$ 25 mm height. The NIR cameras are monochrome and sensitive in both the visible spectrum and NIR, following the sensitivity curve of the CMOS sensors. The three cameras were configured to operate at 30 Hz and $1280\times 720$ resolution, and depth images are obtained by combining their three image channels. See Fig. 1 for more details on the position of each camera.
  • Smartwatches: 2 different smartwatches, the Huawei Watch 2 and the FitBit Sense, were used to monitor heart rate [25, 26, 14], stress level (EDA sensor) [25, 27, 28], inertial sensors (gyroscope and accelerometer) [29]. It was decided to use 2 different smartwatch models, each worn on a different wrist, so that the heart rate monitored by the devices could be compared. The heart rate measurement was set to 1 Hz for the Huawei Watch 2 and 0.2 Hz for the FitBit Sense. The FitBit Sense has one sensors that are not available in the Huawei Watch 2: the EDA sensor. The Huawei Watch 2 was worn on the dominant hand of the learner to detect the movements produced by mouse usage.
  • EEG: A NeuroSky EEG headsetdddhttps://store.neurosky.com/pages/mindwave, which measures the power spectrum density across 5 electroencephalographic channels ($\alpha,\beta,\gamma,\delta,\theta$). Through the preprocessing of these EEG channels, estimates of attention and meditation levels, as well as the occurrence of eye blinks, are obtained. EEG captures voltage signals typically generated by synaptic excitations of the dendrites in the pyramidal cells located in the top layer of the brain cortex [30]. These signals are primarily produced by the synchronous firing of numerous neurons and fibers [31]. EEG is an important, effective, and objective measure for estimating cognitive load and learner states [32, 33] because these signals are sensitive to the mental effort/cognitive works and mental states such as learning, lying, perception and stress. Extensive research in neurology and learning fields has explored the relationship between EEG signals and mental activities [23, 34, 31, 35, 36, 37], demonstrating the potential of the five electroencephalographic channels and the correlation between eyeblinks and cognitive load [38, 21, 23, 39].
  • Visual attention: A Tobii Pro Fusioneeehttps://go.tobii.com/tobii-pro-fusion-product-description equipped with two high-speed infrared cameras configured at 120 Hz for eye tracking. This device estimated the following data: gaze origin and point, pupil diameter, eye movement type (fixation, saccade, unclassified, eyes not found), event duration, data quality, eyeblink, and more; allowing us to measure visual attention. Previous studies have demonstrated that visual attention is a key factor in learner learning, providing a deeper understanding of learner performance [40, 41]. There is a strong correlation between learner performance and visual attention, with features such as fixations, pupil diameter, eyeblink and saccades being closely linked to learning outcomes [40], fatigue detection [42], and task prediction [43].
    Figure 3: Overview of the platforms and tools used to capture the IMPROVE database. The edBB setup was used to monitor the learner’s biometric and behavioral signals during the online session. The edX platform was utilized for delivering the course content and collecting session metadata. The LOGGE Tool was employed to record additional information, such as medical issues and group assignments, and monitor mobile phone usage. The data were aggregated into the IMPROVE database, which can be analyzed using the M2LADS platform to visualize and synchronize the captured signals and events.
  • Computer Information: Relevant information for the session is obtained from various sources on the learner’s computer. i) Microphone Information: Audio is captured from the computer’s microphone at a frequency of 8000 Hz. The audio is recorded throughout the session, and all learners pronounce an initial phrase, "Sound test for the edBB platform," which can be used as a baseline for future research. ii) Keyboard and Mouse Activity: Keystrokes, mouse position, time between keystrokes, mouse wheel movements, etc., are monitored. Keystrokes and mouse dynamics have proven to be effective mechanisms for understanding learner sessions and providing secure certification [44, 45]. iii) Screen Capture: The monitor screen is captured at a frequency of 1 Hz. iv) Session metadata: Information such as keyboard type, logging data, IP and MAC addresses, operating system, etc., is collected.
The edBB platform facilitated optimal synchronization among all sensors, recording the captured samples in both UNIX time and local time, and sending the information to a common server for easy access. By using a single platform to control all sensors, it was possible to synchronize the activation and termination times of each sensor. This also allowed us to obtain precise information, such as the exact moment a learner presses a key and the corresponding video frame, without the need for additional processing.
In addition to using these platforms, a log recording tool called LOGGE [22] was designed, with the primary objective of labeling the mobile phone events of Group 1 and Group 2, which consisted of learner interactions with their mobile phones or when their phones rang. Additionally, this tool was responsible for collecting information related to the monitored learner (e.g., use of glasses/contact lenses, cardiac conditions, handedness, etc.) and enhancing the edX log traces with information related to the activities the learner engages in during the learning session (start and finish times of each task, error logs during the session, etc.).
Finally, a web-based system called M2LADS [22] was designed to visualize the information captured from each monitored learner. This system allowed for the following: i) Validate that the information was correctly monitored and synchronized: visualizing the information from the captured signals (webcam videos, monitor video, heart rate, attention level, etc.) and ensuring their correct synchronization. ii) Statistical analysis of the data: comparing learner performance values (pretest score vs. posttest score), analyzing signals based on activities (identifying which activities had higher attention levels or heart rates, etc.), and correlating biometric signals. The overall architecture, including the platforms and tools used to gather IMPROVE database, is summarized in Fig. 3.
Data Processing
The videos from the RealSense RGB camera were processed to obtain the learner’s facial position, facial landmarks and 3D head posture for each frame. Three state-of-the-art methods based on Convolutional Neural Networks (CNNs) were utilized:
  • Facial Detection: The facial detection module employed MediaPipe’s BlazeFace[46] model (full range) for 2D facial image detection. This module was based on CNNs specifically designed to significantly reduce computational cost and inference time without sacrificing accuracy. BlazeFace integrates a lightweight feature extraction architecture inspired by MobileNetV1/V2[47], leveraging its efficiency and speed. Specifically, the model’s architecture utilizes advanced convolution techniques, similar to those employed in networks like CenterNet[48], with a customized encoder. The detector was trained on large, private datasets that include geographically diverse data based on facial detection.
  • Head Pose Detection: A real-time head pose detector based on the WHENet[49] architecture was employed. The Euler angles —vertical (pitch), horizontal (yaw), and longitudinal (roll)— were estimated through a 2D image of the head, allowing the 3D orientation of the head to be deduced from 2D facial images. This architecture was based on fully connected CNNs for angle classification. Specifically, EfficientNet[50] was utilized as the backbone for feature extraction. Several layers comprised this architecture, starting with convolutional layers that capture spatial features of the image. These convolutional layers were followed by pooling layers. A series of fully connected dense layers were used to interpret the extracted features and perform the classification and regression of the Euler angles. In the classification process, each of the angles (yaw, pitch, and roll) was classified into bins, while regression was used to accurately predict the Euler angle values within those bins. The head pose detector was trained on the 300W-LP[51] and CMU[52].
  • Facial Landmark Detection: The MediaPipe library was used, which estimates 478 3D facial landmarks in real-time from facial images. This enables the extraction of relevant information from facial areas such as the eyes, forehead, mouth, and more. The model is based on a CNN, similar to MobileNetV2[47], but with custom blocks designed for real-time performance. The model was evaluated using a private dataset consisting of 1700 examples uniformly distributed across 17 geographical subregions. The mean error was 2.62, with similar error rates observed for each subregion. Consequently, each video is accompanied by the 478 3D facial landmarks for each frame, along with detailed information about the estimation of each landmark, including its visibility and presence in the image.
The data obtained from the eye tracker was processed to extract relevant features such as fixations and saccades. On average, Tobii Pro Fusion captures a sample with gaze coordinates every 8.3 ms. These samples can be processed for the entire video to obtain relevant visual attention features, such as fixations and saccades. For this purpose, one of the filters integrated into the Tobii Pro Lab software was selected, specifically the I-VT (Fixation) filter. We selected this filter because it is suitable for our study, where head movements are minimal, and the predominant eye movements are fixations and saccades. This filter assigns fixations or saccades to gaze points based on their speed. If the speed is below 30 degrees per second, the points are assigned to the same fixation. Some characteristics of this filter include: i) Noise reduction: A moving average is applied every 3 samples. ii) Velocity calculation: Velocity is calculated in 20 ms windows. iii) Minimum fixation duration: Fixations shorter than 60 ms are discarded. iv) Merging adjacent fixations: Adjacent fixations are merged when two fixations occur with a time interval of 75 ms or less, and the angular distance between them is 0.5 degrees or less.
Figure 4: Example of filtered signals in IMPROVE: The graph on the left shows the raw heart rate signal and the same signal filtered using a moving average filter (with a 15-second window), while the graph on the right presents the raw attention signal obtained from the EEG and filtered using a median filter (with a 5-second window). The original signals are represented by a solid blue line, while the filtered signals are shown with a solid yellow line.
The data obtained from the EEG headband were filtered to mitigate the impact of potential errors, such as signal fluctuations caused by sudden movements of the participant, ensuring more accurate recordings of the EEG signals. The signals were processed using a median filter with a window size of 5, effectively reducing noise caused by artifacts, such as participant movements, eye blinks, and other factors, without distorting the peaks or extreme values in the signal. Additionally, interpolation was applied for data losses shorter than 5 seconds, while longer data gaps were marked as missing information. Both the raw and filtered data are available in the database.
The heart rate data was also filtered to eliminate minor fluctuations and smooth the signal. Specifically, a moving average filter with a 15-second window was applied, as the pulse signal in this context is expected to change slowly, with variations between beats typically not being abrupt at the sampling rate of smartwatches. In this case, only one user experienced a data loss of 60 seconds, which was marked as missing information. The signal obtained from the smartwatch’s accelerometer was filtered using a fourth-order Butterworth low-pass filter with a cutoff frequency of 15 Hz. This was done to eliminate high-frequency noise, such as vibrations or unwanted movements, while preserving the rapid and relevant movements during the learner’s interactions with the mouse. For the gyroscope signal, a similar Butterworth low-pass filter was applied, but with a cutoff frequency of 10 Hz, which allowed for the capture of rapid wrist movements while eliminating potential high-frequency artifacts. Both the raw and filtered data are available in the database. Fig. 4 presents examples of filtered signals from the IMPROVE database.
Data Records
The IMPROVE dataset is available through Science Data Bank[53] (https://doi.org/10.57760/sciencedb.14565) and GitHub[54] (https://github.com/BiDAlab/IMPROVE). The database consists of 2.83 terabytes of information, including biometric data, physiological signals, videos, supervisor-labeled events related to mobile phone usage, and more. This information has been collected from the edBB and edX platforms, as well as from the LOGGE tool, and also includes processed signals. The dataset is organized into three main folders, each corresponding to a different group based on mobile phone usage (see Subsection Protocol for more details). Within each of these folders, there are 40 main subfolders, each labeled with the learner’s ID, containing the monitored data for each learner during the learning session. Each of these folders is further divided into four categories: ’edBB’, ’edBB Processed Data’, ’edX,’ and ’LOGGE,’ corresponding to the information collected from each of the platforms and tools. Tables summarizing the dataset structure and file descriptions—Table 6, Table 7, and Table 8—are provided in the Appendix section for reference. The processed signals and videos obtained through the edBB platform are stored in the ’edBB Processed Data’ folder and are referenced in Table 6. Table 7 lists the files and descriptions from the edX platform and the LOGGE tool. Additionally, all files from the edBB platform and their descriptions are summarized in Table 8.
Technical Validation
For the technical validation of the data, we present a se-supervised data curation method and $3$ different statistical analysis performed using the data over the different groups monitorized.
i) Semi-supervised Event Data Curation:
A semi-supervised method was employed for the data curation of the mobile phone usage events. To this end, we used the eye tracker data and head pose estimation through video (see Fig. 5) for an automatic detection of mobile phone usage events. This automatically annotated events were validated with the events labeled by the human supervisor, who manually labeled the start and end of each event.
Figure 5: Diagram of the proposed semi-supervised method for curating mobile phone usage event data.
The semi-supervised data curation method processes video data from the RGB webcam to detect head pose changes and moments when the learner was not looking at the computer, which with high probability, could correspond to mobile phone usage (see Fig.5). To process the videos and detect head pose changes, we used two state-of-the-art modules based on CNNs, which are described in more detail in the Data Processing section.
  • In the first module (I), the video frames were processed to detect the learner’s face by selecting the most centered and widest bounding box for each frame, using MediaPipe’s BlazeFace facial detector [46].
  • In the second module (II), the WHENet head pose estimator [49] was applied to obtain the Euler angles (pitch, roll, and yaw).
  • In the third module (III), the method presented by Becerra et al. [55] was followed to detect abnormal head pose events. The process involves calculating the Euler angles (pitch, yaw, and roll) for each video frame. A sliding window ($W_{l}$) methodology was then applied to compute the average angles within each temporal window. An event was detected when the local average of an angle in a window exceeded a predefined threshold, which was determined by a significant deviation from the global average. This method ensures that only meaningful head pose changes are flagged as events, minimizing false positives from minor movements. The method consists of the following steps:
    1. 1.
      The mean ($\mu$) and standard deviation ($\sigma$) for each Euler angle were calculated for all sessions.
    2. 2.
      For each temporal window $W_{l}$, the average $\mu$ of the Euler angles was calculated, providing a representative value for pitch, yaw, and roll within that specific time frame.
    3. 3.
      A temporal window $W_{l}$ was labeled as an event when the $(\mu_{W_{l}})$ exceeds a predefined threshold, which was set based on a significant deviation from the global mean ($\mu$) of the Euler angles calculated over all sessions. The following equation shows the threshold: $\left|\mu_{W_{l}}^{i}-\mu^{i}\right|>n\sigma^{i}$, where $i$ represents the angle (yaw, pitch, or roll) and $n$ is the threshold factor that defines the sensitivity of the event detection, determining how many standard deviations away from the mean a particular value must be to be considered an event.
    4. 4.
      All overlapping events were merged.
    Figure 6 shows the detection of abnormal head pose events applied in module c), for two learners from the IMPROVE dataset.
  • In the fourth module (IV), information from the eye tracker was used to discard abnormal head pose false detections in cases where the learner continued to look at the computer screen.
  • In the fifth module (V), the information from the LOGGE tool was used to determine when mobile phone usage events were labeled. This information helped select only the predicted events from our abnormal head pose detector that occurred within the time window $W_{s}$ indicated by the supervisors, including a buffer $\Delta t=10\,\text{s}$ before and after the event. The assumption is that manual labeling was generally accurate but involved slight delays when marking the start and end of each event. The conditions for maintaining the supervisor’s labeling were as follows: the abnormal head pose detector must report the same number of events within $W_{s}\pm\Delta t$ as those recorded by LOGGE, and the overlap between the predicted and manually labeled events must have a delay of less than 4 seconds. If these conditions are not met, manual re-labeling is required using the TagJump application we developed. TagJump allows displaying the video of the session at the start and end times of events reported by LOGGE and those predicted by the abnormal head pose detector. It also enables quick navigation through nearby frames to facilitate re-labeling. The protocol of this application first shows the LOGGE-labeled event, and if the event is not found within the first 4 seconds, the event predicted by the detector is displayed. This approach simplifies the supervisor’s labeling process.
Figure 6: Example of the abnormal head pose detector for two students. The graphs show the Euler angles obtained during the session, along with the average for each angle in a window $W_{l}=4s$. The set thresholds for each angle are also shown, and when these are exceeded, they are labeled as events. The detector exhibits a high overlap with the manually labeled events. The images show the students’ pose while using a mobile phone. Yaw is represented by the solid blue line, and roll by the solid red line. The temporal window average is indicated by the dashed yellow line. Predicted event labels are framed in red, while manually labeled events are framed in green. Thresholds are represented by dotted lines with the corresponding color for each angle.
Figure 7: Average event detection sensitivity obtained by the detect abnormal head pose approach to detect mobile phone usage across all users (left). Average number of detected events by the detect abnormal head pose approach across all users within the window $W_{s}\pm\Delta t$ (right). The x and y axes in both heatmaps represent the tested parameters $W_{l}$ and $n$, respectively, which were varied to analyze their effect on the detection performance.
To fine-tune the parameters $W_{l}$ and $n$, a preliminary study was conducted using data labeled by the LOGGE tool (see the previous subsection Data, Platforms, and Sensors), specifically focusing on the controlled events, which were presumed to have been labeled with greater precision by the supervisors. The goal was to optimize the parameters to maximize the detection of mobile phone usage events while minimizing false positives, ensuring that the predicted windows closely matched those labeled by the supervisors. Fig. 7 presents two heatmaps for controlled events. The first heatmap illustrates the event detection accuracy, while the second shows the average number of detected events within the windows $W_{s}\pm\Delta t$. Both are dependent on the selected parameters and the temporal window $W_{s}\pm\Delta t$. Based on the data analysis, we selected $W_{l}=4s$ and $n=2$ as the optimal parameters. However, this choice is not straightforward, and other combinations of parameters could also be valid depending on the specific use case.
Mean (Std)
Sensors Variable (Unit) Group 1 Group 2 Group 3 Overall
Smartwatch Heart Rate (bpm) 82.85 (10.25) 79.82 (13.70) 78.26 (14.30) 80.31 (12.75)
Accx (m/s2) 0.98 (0.48) 0.81 (0.58) 0.75 (0.50) 0.85 (0.52)
Accy (m/s2) 5.05 (1.39) 4.19 (1.54) 4.82 (1.39) 4.69 (1.44)
Accz (m/s2) 7.51 (0.89) 8.46 (0.94) 8.13 (0.95) 8.03 (0.92)
Gyrox (°/s) 0.07 (0.03) 0.04 (0.03) 0.04 (0.02) 0.05 (0.03)
Gyroy (°/s) 0.04 (0.01) 0.02 (0.01) 0.02 (0.01) 0.03 (0.01)
Gyroz (°/s) 0.04 (0.02) 0.03 (0.02) 0.03 (0.01) 0.03 (0.02)
EEG Band Attention (0-100) 50.91 (5.68) 54.98 (4.40) 51.21 (6.54) 52.37 (5.54)
Meditation (0-100) 56.39 (4.46) 54.25 (5.48) 54.58 (5.79) 55.07 (5.24)
Alpha (dB) 6.38 (0.62) 5.47 (0.61) 6.23 (0.61) 6.03 (0.61)
Beta (dB) 3.32 (0.62) 3.28 (0.58) 3.67 (0.58) 3.42 (0.59)
Delta (dB) 1.70 (0.97) 1.59 (1.00) 1.59 (0.95) 1.63 (0.97)
Gamma (dB) 1.89 (0.90) 1.97 (0.82) 2.06 (0.75) 1.97 (0.82)
Theta (dB) 10.02 (0.69) 9.21 (0.71) 9.84 (0.71) 9.69 (0.70)
Table 4: Absolute mean and standard deviation for physiological and biometric signals across the different groups in the IMPROVE dataset.
Following this protocol, 80% of the LOGGE events were re-labeled. Only 20% had a delay of less than 4 seconds. The LOGGE labels and those detected by the abnormal pose detector complemented each other. In 24.80% of the events, when one of the labels had a delay of more than 4 seconds, the other label had a delay of less than 4 seconds. This complementarity between the two labeling systems facilitated faster re-labeling with the TagJump application. 27.2% of the re-labeled events showed a difference of less than 3 seconds, 15.2% had a difference between 3 and 4 seconds, while 57.6% showed a delay of more than 4 seconds compared to the LOGGE reports.
The re-labeled events were checked to determine whether they improved the event detection accuracy of the abnormal head pose detector. For controlled events, accuracy improved by approximately 3%, and for uncontrolled events, by around 5%, demonstrating that the detector becomes even more accurate with proper re-labeling.
Figure 8: Comparison of heart rate measurements between the Huawei 2 and Fitbit Sense for two different students. The solid blue line represents the Huawei 2, while the orange dashed line corresponds to the Fitbit Sense.
ii) Statistical Distributions of Physiological and Biometric Signals:
The mean and standard deviation of each biometric and physiological signal were calculated to ensure that no general errors were present and that the values fell within expected ranges for a similar environment (see Table 4). Additionally, these results were compared with the mEBAL2 dataset [21], which was obtained in a similar e-learning environment and utilized the same EEG band. The values obtained for the various physiological and biometric signals are found to be within the expected ranges for learners in an e-learning context. The average heart rate across the three groups ranged from 78.26 to 82.85 bpm, consistent with values reported in previous studies on students during virtual learning sessions [26], [14], and with the normal human resting heart rate, typically between 60-100 bpm [56].
The inertial sensors (accelerometer and gyroscope) in the Huawei Watch 2 smartwatch, worn on the hand used to operate the mouse, provided the following observations. The average acceleration in the X and Y axes reflects low values, which is expected during a learning session where most of the time is characterized by either no movement or minor mouse movements. In contrast, the higher values on the Z axis are primarily attributed to the influence of gravitational force. Similar results are observed in the gyroscope, with mean values close to zero across all axes, reflecting the same reasons.
For attention and meditation, the means were 52.37 and 55.07, respectively, aligning with previous studies in e-learning environments, such as the mEBAL2 dataset [21, 8], which also report values around 50% for attention and 53.88% for meditation. These findings are in line with moderate-length sessions, where attention generally starts high but gradually decreases over time, stabilizing around the mean without extended periods of low attention [11]. EEG signals (Alpha, Beta, Delta, Gamma, Theta): The average values of the EEG frequency bands in IMPROVE are consistent and comparable to those obtained in the mEBAL2 dataset, where Alpha (8.09 dB), Beta (5.22 dB), Delta (2.00 dB), Gamma (2.59 dB), and Theta (9.93 dB) are reported. In both datasets, the order of the signals according to their relative power remains constant, and the value ranges are very similar. The slight differences observed can be attributed to the varying learning activities performed in each dataset, as well as the intrinsic differences between users.
iii) Heart Rate Data Interoperability between Devices:
The heart rate data from both smartwatches (Huawei Watch 2 and FitBit Sense) were compared to ensure consistent measurements, as similar values would indicate accurate data recording. The heart rate sampling frequency is different between the two devices, with data being captured by the Huawei Watch 2 at 1 Hz, while the FitBit Sense operates at a lower frequency of 0.2 Hz. A difference of 3.70 bpm was found between the two signals (see Fig. 8), which falls within the reasonable range for this type of device, typically around ±$\pm 3\text{-}5$ bpm [57]. Additionally, it should be noted that the smartwatches were worn on different wrists, with the dominant hand experiencing more movement, potentially affecting heart rate estimation. This, along with the variation in sampling frequencies, may have contributed to the observed differences.
Mean (Std) Statistical Tests
Sensors Variable (Unit) No Phone Phone All Sessions t p-value
Smartwatch Heart Rate (bpm) 82.68 (10.3) 84.03 (10.29) 82.85 (10.25) -2.56 0.0153
Accx (m/s2) 0.74 (0.48) 2.52 (0.99) 0.98 (0.48) -10.82 0.0000
Accy (m/s2) 4.58 (1.55) 8.05 (1.03) 5.05 (1.39) -12.15 0.0000
Accz (m/s2) 8.17 (1.00) 3.31 (1.25) 7.51 (0.89) 16.80 0.0000
Gyrox (°/s) 0.05 (0.02) 0.22 (0.09) 0.07 (0.03) -11.25 0.0000
Gyroy (°/s) 0.03 (0.01) 0.09 (0.04) 0.04 (0.01) -10.50 0.0000
Gyroz (°/s) 0.04 (0.01) 0.09 (0.04) 0.04 (0.02) -9.70 0.0000
EEG Band Attention (0-100) 51.28 (5.5) 48.47 (8.05) 50.91 (5.68) 3.03 0.0048
Meditation (0-100) 56.33 (4.4) 56.80 (6.51) 56.39 (4.46) -0.59 0.5626
Alpha (dB) 6.31 (1.72) 6.85 (1.78) 6.38 (1.69) -3.57 0.0011
Beta (dB) 3.30 (0.95) 3.45 (0.98) 3.32 (0.94) -2.00 0.0541
Delta (dB) 1.69 (0.43) 1.80 (0.49) 1.70 (0.43) -2.35 0.0250
Gamma (dB) 1.86 (0.80) 2.03 (0.72) 1.89 (0.78) -2.36 0.0243
Theta (dB) 9.99 (2.36) 10.18 (2.13) 10.02 (2.31) -0.88 0.3839
Table 5: Absolute mean and standard deviation for physiological, biometric, and movement signals across phone and no phone conditions, with t-values and p-values highlighting statistical significance, for Group 1 in the IMPROVE dataset.
iv) Statistical Validation of Signal Changes Due to Mobile Phone Usage:
The absolute mean and standard deviation of the biometric and physiological signals were calculated both during mobile phone usage events and when the mobile phone was not in use. A statistical analysis was then conducted to determine whether significant changes occurred in the signals during mobile phone usage. The results in Table 5 indicate that, according to the Student’s t-test, almost all signals have p-values below 0.05, suggesting a low likelihood that the observed differences between phone and no-phone conditions are due to chance. Overall, these findings suggest that significant impacts on these variables are caused by mobile phone usage.
Notably, the accelerometer and gyroscope signals exhibit highly significant differences between the phone and no-phone conditions, as expected due to the movement involved when using a mobile phone, which contrasts with typical movements during an online session. Heart rate also shows a significant difference, indicating that the heart rate is affected by mobile phone use. Most EEG signals, including attention, also show significant differences. "However, the Meditation, Beta, and Theta signals do not exhibit statistically significant differences with respect to the mean. Theta waves are related to relaxation and wakeful states, so it is unsurprising that the Meditation signal shows no significant differences, suggesting that relaxation or meditation is not notably affected by mobile phone usage. However, further research is required to confirm this.
Beta waves are associated with focus, normal waking consciousness, and alertness [35, 36]. For instance, Beta waves occur during problem-solving, decision-making, and anxious thinking. Nevertheless, Beta did not show statistically significant differences in this study.
Figure 9 presents boxplots for attention and Theta waves across two experimental conditions: "Without Using Phone" and "Using Phone." Attention levels tend to be lower when the mobile phone is used, while the range of values is wider when the phone is not in use. In contrast, Theta waves show a slight increase during mobile phone usage.
In summary, most signals are affected by mobile phone usage, validating the correct labeling of these events and supporting the potential to study the effects of mobile phone use through the IMPROVE dataset.
iv) Verification of Signals and Video Recordings with M2LADS
M2LADS [22] was used to validate the signals and video recordings captured during the session. Through the system’s visualization and the synchronization it provides, all data were checked, and no corrupt files were reported by the system. The results of the validation process using M2LADS were as follows:
  • EEG Band: 35% of the learners experienced some data loss, with 23.3% of users losing 1 minute of data and 14.16% losing up to 5 minutes of data. These losses were primarily caused by abrupt head movements, which dislodged the EEG band, occurring more frequently among female participants. Supervisors closely monitored the sessions to prevent such issues.
  • Depth Video: One depth video file was found to be corrupt for the user with ID: 202303281.
  • Left NIR Camera: A corrupt file was identified for the left NIR camera in the session for the user with ID: 202305233.
  • Frontal RGB Camera: It was reported that an RGB camera video (ID: 202307122) continued recording a few minutes after the session had ended, instead of stopping at the session’s conclusion.
  • Fitbit Sense: The Fitbit Sense data were lost for three students (IDs: 202303272, 202303273, and 202303274).
Figure 9: Boxplot comparing mobile phone usage and non-usage conditions for attention levels (left) and theta wave activity (right).
Usage Notes
Potential applications. The IMPROVE dataset offers a unique opportunity for researchers across various disciplines. Its multimodal data, including physiological signals, video recordings, learning analytics, and learner feedback, provide numerous potential applications. The primary goal of the IMPROVE dataset is to assess the effects of mobile phone usage on learners during online learning sessions. IMPROVE serves as a valuable resource to explore how interaction with mobile phones affects cognitive performance and learner well-being.
Machine learning and signal processing researchers can leverage this dataset to develop and refine models that detect the impact of mobile phone usage on learning performance. The combination of different data sources, such as EEG bands, inertial signals, gaze tracking, heart rate monitors, and video from cameras, enables the exploration of complex patterns that could lead to enhanced algorithms for estimating attention, cognitive load, and the distractions caused by mobile phone usage.
This dataset is also valuable for educational researchers studying the interaction between physiological responses and online learning environments. By analyzing both biometric signals and academic performance data, researchers can investigate the effects of mobile phone usage on cognitive performance, attention span, and learning outcomes.
Similar to the edBB dataset [14, 1], IMPROVE opens up challenges to the video processing community, including tasks related to RGB, NIR, and depth camera data, such as pulse estimation, blink detection, user pose estimation, and detection of abnormal learner behavior, among others.
Moreover, researchers in Human-Computer Interaction (HCI) could utilize the IMPROVE dataset to develop adaptive learning systems that respond to the physiological states of learners [1, 22].
Psychology and neuroscience communities can also benefit from this dataset by investigating mobile phone use’s cognitive and emotional effects in educational settings. EEG signals, particularly those related to attention and relaxation (e.g., Theta and Beta waves), can be analyzed to explore how cognitive processes are impacted during online sessions involving varying levels of mobile phone interaction.
In summary, the IMPROVE dataset offers extensive potential for applications in educational technology, machine learning, psychology, and human-computer interaction, making it a valuable resource for interdisciplinary research into the effects of mobile phone usage in digital learning environments.
Suggestions: It is highly recommended to use the filtered EEG band files or apply filters using programs like Matlab or Python. In Python, the SciPy library (https://scipy.org) provides algorithms for signal processing, such as signal filtering. The EEG band is often affected by artifacts due to movement or eyeblinks, so it is recommended to apply a moving average or median filter to improve the quality of the signals [8].
The Pandas library (https://pandas.pydata.org/) proved very useful for manipulating and analyzing large datasets, while NumPy (https://numpy.org) was used for processing any type of signal or video. For training deep learning models or neural networks, TensorFlow (https://www.tensorflow.org/) and PyTorch (https://pytorch.org/) are two of the most powerful and flexible platforms for this purpose, and were essential for running the models applied in video processing.
For video processing, the Dlib library (www.dlib.net), which includes pre-trained models for facial detection, facial landmark estimation, and more, was used. This library integrates well with OpenCV (https://opencv.org/), which facilitates video reading, filtering, distortion correction, and also provides pre-trained models. This combination was used to process the videos, alongside MediaPipe (https://github.com/google-ai-edge/mediapipe?tab=readme-ov-file), a library specializing in fast models for face detection, facial landmarks, image segmentation, and more. MediaPipe was employed to extract the facial bounding boxes and landmarks available in the IMPROVE dataset. The use of RetinaFace (https://github.com/serengil/retinaface) is also recommended for facial detection with facial landmarks.
For head pose estimation, WHENet (www.github.com/Ascend-Research/HeadPoseEstimation-WHENet) is a fast and efficient solution, which was used to obtain the Euler angle labels in IMPROVE. Another efficient option for head pose estimation is RealHePoNet (https://github.com/rafabs97/headpose_final).
For processing eye-tracker data, the official software Tobii Pro Lab (https://www.tobii.com/products/software/behavior-research-software/tobii-pro-lab) is recommended, which was used in IMPROVE. Another interesting solution is PsychoPy (https://www.psychopy.org/), a free tool offering support for the creation and execution of experiments involving eye tracking.
Data Access: The IMPROVE dataset is available exclusively for research purposes and must not be used or included in any commercial applications in any form (e.g., original files, encrypted files, files containing extracted features, etc.). Researchers wishing to access the dataset can do so through two available methods. The license agreement can be downloaded from GitHub (https://github.com/BiDAlab/IMPROVE) and a signed, scanned copy should be sent via email to atvs@uam.es, following the instructions provided on the GitHub repository. Alternatively, access can be requested through the Science Data Bank (https://doi.org/10.57760/sciencedb.14565), where the necessary steps for obtaining the data are outlined on the platform.
Code Availability
The code for filtering physiological and biometric signals, as well as detecting pose events, is available on GitHub (https://github.com/BiDAlab/IMPROVE). MATLAB was used for signal filtering, while Python was utilized for pose event detection. All necessary packages are listed in the requirements.txt file. These available codes aim to assist the community in processing the IMPROVE dataset for various research and analysis purposes.
Acknowledgements
This work has been supported by projects: HumanCAIC (TED2021-131787B-I00 MICINN), BIO-PROCTORING (GNOSS Program, Agreement Ministerio de Defensa-UAM-FUAM dated 29-03-2022), SNOLA (RED2022-134284-T), e-Madrid-CM (S2018/TCS4307). Research partially funded by the Autonomous Community of Madrid. Roberto Daza is supported by a FPI fellowship from MINECO/FEDER. A. Morales is supported by the Madrid Government (Comunidad de Madrid-Spain) under the Multiannual Agreement with Universidad Autónoma de Madrid in the line of Excellence for the University Teaching Staff in the context of the V PRICIT (Regional Programme of Research and Technological Innovation).
Author Contributions Statement
R.D.: Conceptualization, Data Collection, Data Curation, Data Analysis, Investigation, Software, Technical Validation, Writing—Original Draft, Writing—Review & Editing. A.B.: Data Collection, Data Curation, Software, Data Analysis, Investigation, Writing—Review & Editing. R.C.: Funding Acquisition, Investigation, Supervision, Writing—Review & Editing. J.F.: Funding Acquisition, Supervision, Writing—Review & Editing. A.M.: Project Administration, Funding Acquisition, Supervision, Writing—Review & Editing.
Competing interests
The authors declare no competing interests.
Appendix: Comprehensive Overview of Data Collected and Processed in the edBB, LOGGE, and edX Platforms
Folder Name Files Description Format File Contents
Video box
head_pose
landmarks
Processed video data: Bounding boxes, facial landmarks, and head pose of all users in the RealSense video .csv
.csv
.csv
Bounding box is formatted as [x, y, width, height], 478 landmarks per detected face, Euler angles (yaw, pitch, and roll) in degrees, frames
MindWave filter_ATT
filter_MED
filter_BANDPOWER
EEG band signals filtered with a median filter .csv
.csv
.csv
Attention (0-100), meditation (0-100), blink strength, 5 EEG waves (dB), local date and time
Smartwatch filter_ACC
filter_Gyro
filter_Heart
Huawei Watch 2 signals filtered using a Butterworth low-pass filter or a moving average filter .txt
.txt
.txt
Acceleration on X, Y, Z axes (m/s2), angular velocity on X, Y, Z axes (${}^{\circ}/s$), heart rate (bpm), UNIX timestamp
Fitbit Heart Heart rate signal from the FitBit Sense filtered with a Butterworth low-pass filter .csv Heart rate (bpm), timestamp
Table 6: Description of Processed edBB Files, Contained in the edBB Processed Data Folder.
Folder Name Files Description Format File Contents
Logge file_events_LS
file_marks
file_position
user_data
file_prestest
Activities and events logs .csv
.csv
.csv
.csv
.csv
Key name for the activity/event, start timestamp, end timestamp, marks of assignments, mobile phone position in table, medical issues, group, pretest mark, posttest mark, self-reported stress before and after the session, self-reported distraction before and after the session, self-reported difficulty of session, self-reported performance
edX file edX logs .csv Action name, context, session, time, event type
Table 7: Description of Files Collected by the edX Platform and LOGGE Tool, Contained in the edX and LOGGE Folders.
Folder Name Files Description Format File Contents
MouseCapture Mouse_Event
Mouse_Move
Mouse_Dragged
Mouse_Wheel
Mouse_Click
Mouse click events and movement tracking .csv
.csv
.csv
.csv
.csv
Press/Release events, mouse button pressed, position (pixels), drag-and-drop detection, scroll movement, click count, Unix timestamp, frame from the webcams
Keyboard Keyboard_Capture
Keyboard_Key_Typed
Keyboard data capture and typed Unicode characters .csv
.csv
Press/Release events, key code, native key code, associated native Unicode character, Unix timestamp, frame from the webcams
MindWave file_ATT
file_MED
file_BLINK
file_BANDPOWER
EEG data .csv
.csv
.csv
.csv
Attention (0-100), meditation (0-100), blink strength, 5 EEG waves (dB), local date and time
Smartwatch ACC
Gyro
Heart
Huawei Watch 2 data .txt
.txt
.txt
Acceleration on X, Y, Z axes (m/s2), angular velocity on X, Y, Z axes (${}^{\circ}/s$), heart rate (bpm), UNIX timestamp
Fitbit Heart
EDA
FitBit Sense data .csv
.csv
Skin conductance levels, heart rate (bpm), timestamp
Eye Tracker Learner ID All information obtained from the eye tracker .tsv Recording timestamp (µs), recording date (local date of the recording), recording start time (local start time), recording resolution (pixels), gaze point (pixels), pupil diameter (mm), eye movement type, validity
VideoMonitor monitor
Videomonitor
Screen recording .csv
.mp4
Frame and timestamps recorded in both UNIX and local time
VideoWebcam webcam
WebcamCapture
Side webcam recording .csv
.mp4
Frame number and timestamps recorded in both UNIX and local time
VideoWebcam2 webcam
WebcamCapture
Overhead webcam recording .csv
.mp4
Frame number and timestamps recorded in both UNIX and local time
RealSense Time
Color
Depth
Left_Infrared
Right_Infrared
Front-facing RGB, NIR, and depth cameras .csv
.mp4
.mp4
.mp4
.mp4
Frame number and timestamps recorded in both UNIX and local time for each camera
SoundCapture Record Session audio recording .wav
PCInformation PCInformationCapture PC information .csv Computer Name, IP address, MAC address, operating system and version, system architecture, main HDD, free HDD, screen resolution, keyboard language
StudentData StudentData Session and learner information .csv Learner ID, session number, session start/end time (local time), session completion status indicator
Table 8: Description of Files Collected by the edBB Platform and Contained in the edBB Folder.